# Dynamics in an emergent quantum-like state space generated by a nonlinear classical network

### v1

@@ IV.1 Phase and the adjacency matrix
This unitary transformation of $A$ does not change the spectrum, but it does rotate the basis for the eigenstates. We use this transformation below, after defining projection of the emergent states, to show how the eigenvectors of the projected emergent states transform under continuous rotations. Later in this paper, we use the transformation of Eq. 13 to account for how oscillator phases in the classical network dictate the eigenvector basis. That basis change, in general, varies throughout the ensemble and therefore influences the mixedness of the state[53]. Recall that we use the term ‘ensemble’ to mean an average over initial phases.
Notice that the matrix $\Phi$ lists the phase of each oscillator in the system, associating each phase to a vertex representing an oscillator. Recall that this phase is the sum of accumulated phase associated with the frequency offset from the mean and a phase offset, $\theta_{i}(t)=\epsilon_{i}t+\phi_{i}$. The transformation weights edges in the graph by phase differences. Thus, for an edge connecting vertex $j$ to vertex $k$, the entry in the upper triangular part of the adjacency matrix transforms as $1\rightarrow\exp(-i[\theta_{j}-\theta_{k}])$, with the corresponding complex conjugate map in the lower triangular part. The diagonal elements remain as entries of 0.
In the numerical examples shown below we obtain the phases from simulations using the Kuramoto model, described in the previous section. For any time point during the simulation, we use the oscillator phases to define $\Phi$, then we diagonalize $\Phi^{-1}A\Phi$ to obtain the eigenvectors of the emergent states—now with complex coefficients. We project these coefficients to those for an effective two-state system for each QL bit, as we now describe for the example of a single QL bit. The method is generalized in the obvious way for a system comprising any number of QL bits.
@@ IV.2 The effective 2-state model for QL bit emergent states
An arbitrary emergent state, $W$, of the QL bit graph will be written in terms of this basis as:
$W=c_{1}u_{1}+\dots+c_{n}u_{n}+d_{1}x_{1}+\dots+d_{k}x_{k}$ (14)
where the $c_{i}$ and $d_{j}$ are complex coefficients. Now define
@@
$\alpha=\langle J_{a1},W\rangle$ (17)
$\beta=\langle J_{a2},W\rangle$ (18)
to give
$W_{2\times 2}=\alpha|a_{1}\rangle+\beta|a_{2}\rangle.$ (19)
Some examples of coefficients for the effective two-states of a QL bit are provided in Fig. 2. Each panel shows a vector representation of the complex coefficients of $W_{2\times 2}$ corresponding to the highest eigenvalue ($\alpha_{0}$ is violet red, $\beta_{0}$ is light pink) and the second highest eigenvalue ($\alpha_{1}$ is dark blue, $\beta_{1}$ is light blue). Fig. 2a shows the eigenvectors for partially synchronized oscillators in one QL bit, derived from phases early in the dynamics of a phase oscillator system with initially random phases. The simulation used the same parameters as detailed later for the simulations of synchronization on the product graph of two QL bits. These coefficients contrast to those obtained for the synchronized oscillators in one QL bit, where phases were obtained later in the simulation once the oscillators has synchronized, Fig. 2b.
@@ IV.3 Phase in the states
Figure 2: Vector representation of the complex coefficients of $W_{2\times 2}$ corresponding to the highest eigenvalue ($\alpha_{0}$ is violet red, $\beta_{0}$ is light pink) and the second highest eigenvalue ($\alpha_{1}$ is dark blue, $\beta_{1}$ is light blue). (a) Examples of systems comprising partially synchronized oscillators. (b) Examples of systems where the oscillators are synchronized.
In the latter case, of a synchronized oscillator system on the QL bit graph, we find that $\alpha_{0}=\beta_{0}\approx 1/\sqrt{2}$ and $\alpha_{1}=-\beta_{1}\approx 1/\sqrt{2}$ (we use the convention that $\beta_{1}$ takes the negative sign). More generally, we can write the eigenvectors in column format and note that they can be equivalently written in the general form, with $\alpha,\beta\in\mathbb{C}$:
$\frac{1}{\sqrt{2}}\begin{bmatrix}1&-1\\ 1&1\end{bmatrix}\rightarrow\begin{bmatrix}\alpha&-\beta\\ \bar{\beta}&\bar{\alpha}\end{bmatrix}$ (20)
Thus, the eigenvectors form the $\mathsf{SU}(2)$ group under the continuous rotations produced by possible phases of the oscillators in each of the two subgraphs. Recall that the special case of negative-signed edges connecting the subgraphs can be interpreted as phases of the two subgraphs being perfectly out of phase[1]. We can visualize these property-preserving rotations using a vector on the unit 3-sphere in four dimensions, $\mathbb{S}^{3}$, or on the Bloch sphere by using the normalization condition $|\alpha|^{2}+|\beta|^{2}=1$. The rotations preserve phases of the oscillators relative to each other and generate $\mathsf{SU}(2)$ on the condition that phases are synchronized. Synchronization can mean the oscillators are all locked in-phase, or subgraphs can be perfectly out-of-phase with each other. Perfect synchronization produces states isomorphic to, for example, the states of a system of spins.
@@ IV.4 Synchronization of quantum bits
$i\hbar\dot{U}_{i}U^{\dagger}_{i}=H_{ii}-\frac{iK}{2N}\sum_{j=1}^{N}a_{ij}(U_{j}U^{\dagger}_{i}-U_{i}U^{\dagger}_{j}).$ (21)
@@
We can emulate this synchronization of quantum bits by building a system of QL bits, interacting via edges connecting vertices of one QL bit to another. For consistency with the Lohe model, we would require each QL bit to comprise strongly synchronized oscillators, and to be synchronized. The emergent state of each QL bit gives its frequency. We can assign a random initial ‘phase’ as the vector in $\mathbb{S}^{3}$ that defines the orientation of the eigenvector of the synchronized emergent state. The system of $N$ QL bits would be combined into a new graph $G=G_{A}\oplus G_{B}\oplus\dots\oplus G_{N}$. The subgraphs are connected by edges, weighted so that the inter-QL bit coupling is weaker than the coupling between vertices within a QL bit. The graphs $G_{A},G_{B},\dots$ are the vertices of the Lohe model. This system of oscillators, propagated according to the Kuramoto model, will mimic the Lohe model. To compare the results we should project out the effective 2-state emergent states of each QL bit to construct the vectors $\mathbf{x}_{i}$.
We do not report these numerical calculations here, but provide the strategy for the following reasons. First, we will refer to this model in a later section of the paper. Second, by proposing how a model comprising a system of coupled classical phase oscillators can predict the same results as a system of coupled quantum bits emphasizes how the QL model succeeds in various ways to produce state spaces very much like those enjoyed by quantum systems.
@@ V.1 Methods
Here we study a network phase oscillator model with a coupling structure templated by the Cartesian product of two QL bit graphs $G_{A}\Box G_{B}$. The phase oscillators are indicated by vertices of the graph, and are assigned a phase that biases the edges, as we describe below.
@@
In the numerical simulations, each graph comprises $n_{0}$ vertices, giving a total of $N=n_{0}\times n_{0}$ vertices in the product graph. Each subgraph $G_{a1}$, $G_{a2}$, $G_{b1}$, $G_{b2}$ comprises 20 vertices (therefore $n_{0}=40$) and is $d$-regular with $d=15$. Edges between subgraphs were added randomly, with probability 0.2. We calculate the full Cartesian product, yielding $2d$-regular subgraphs in $G_{A}\Box G_{B}$, with an eigenvalue of 30 for the emergent state when the subgraphs are uncoupled (that is, neglecting eigenvalue shift labeled $\Delta$ in Ref. [2]). We expect four emergent states, two with an eigenvalue of 30, one at $30+\Delta$, and one at $30-\Delta$.
@@
The graph provides the coupling map for a system of $N$ phase oscillators that we model using the Kuramoto model, Eq. 10. Couplings between vertices are weighted uniformly by $K/N$, with $K$ chosen to be 250 for the first set of calculations. Vertices are assigned a mean frequency of 100 in the same arbitrary units as $K$, and offset by frequencies randomly selected from a normal distribution with standard deviation 1 and mean of 0 (in the same arbitrary units). Initial phases for the oscillators are randomly assigned from a von Mises distribution with circular standard deviation of either 50, meaning the distribution is similar to a uniform distribution, or 0.001, meaning the distribution is extremely narrow. Results are averaged over either 100 or 500 realizations of this set-up. See ref [53] for some background and examples that might be helpful. The coupling $K/N=0.5625$ is moderately strong compared to the ratio of the standard deviation of the frequency distribution to its mean, 0.01.
@@
The main motivation for the ensemble averaging over realizations of the system is to account for uncertainty in the initial oscillator phases. Therefore, we use the term ‘ensemble’ to mean an average over initial conditions of oscillator phase. The initial phase for any oscillator means its natural frequency and its phase offset, $\epsilon_{i}$ and $\phi_{i}$, respectively. This setup for the calculations means that the dynamics are not strictly deterministic, because we do not know the initial conditions precisely. We also report calculations where we, additionally, average over an ensemble of graphs. In fact, because the subgraphs are $d$-regular (with no edges randomly deleted like we showed in prior work[1]), the ensemble comprises graphs with different configurations of edges connecting the subgraphs.
@@
During the evolution of the system, the phase of each oscillator $\theta_{i}$ evolves nonlinearly, as discussed in the previous section. At selected time points we transform the adjacency matrix of the graph according to Eqs. 12, 13, which introduces phase differences by weighting the edges between vertices $j$ and $k$ by $e^{i(\theta_{k}-\theta_{j})}$. We average the spectrum of the transformed adjacency matrix over all iterations of the simulation to give ensemble spectra. The spectra were computed for various time points, and therefore order parameters, as the system synchronized.
Before discussing these results, we note that although the phases of the initial distribution of oscillators are randomly distributed, the phase differences contain embedded correlations—there are $N$ phases, but much more than $N$ phase differences. However, the phase differences are uncorrelated from one simulation to another, so mixedness comes from an essential ensemble averaging over simulations. In prior work mixedness as a function of synchronization was studied in some detail[53].
@@
In Eq. 12 we are applying a unitary transformation to the adjacency matrix. Therefore the spectrum does not change with order parameter, only the state. The density matrix for the relevant emergent state is calculated in the basis produced by the effective two-state model for each QL bit. That is, we statistically average over the pure states:
$\Psi(t)=c_{1}(t)|a_{1}\rangle|b_{1}\rangle+c_{2}(t)|a_{2}\rangle|b_{1}\rangle+c_{3}(t)|a_{1}\rangle|b_{2}\rangle c_{4}(t)|a_{2}\rangle|b_{2}\rangle,$ (25)
for each time point in a simulation. Here the $c_{i}(t)$ are complex coefficients obtained from the projections outlined in Sec. IV.B, and the basis states are the emergent states from each subgraph. For example, for subgraph $G_{a1}$ the emergent state is $|a_{1}\rangle$ and for subgraph $G_{a2}$ the emergent state is $|a_{2}\rangle$. From the ensemble of $\Psi(t)$ we determine the density matrix $\rho(t)$ in the product basis, with entries
$\rho_{ij}(t)=\langle c_{i}(t)c^{*}_{j}(t)\rangle$ (26)
where the angle brackets mean an average over the ensemble of simulations. We report $\mathrm{Tr}(\rho)^{2}$ as a simple measure of mixedness of the state.
V.2 Synchronization

### v3

@@ IV.1 Phase and the adjacency matrix
This unitary transformation of $A$ does not change the spectrum, but it does rotate the basis for the eigenstates. We use this transformation below, after defining projection of the emergent states, to show how the eigenvectors of the projected emergent states transform under continuous rotations. Later in this paper, we use the transformation of Eq. 13 to account for how oscillator phases in the classical network dictate the eigenvector basis. That basis change, in general, varies throughout the ensemble and therefore influences the mixedness of the state[40]. Recall that we use the term ‘ensemble’ to mean an average over initial phases to account for state preparation.
The matrix $\Phi$ lists the phase of each oscillator in the system, associating each phase to a vertex representing an oscillator. Recall that this phase is the sum of accumulated phase associated with the frequency offset from the mean and a phase offset, $\theta_{i}(t)=\epsilon_{i}t+\phi_{i}$. The transformation weights edges in the graph by phase differences. Thus, for an edge connecting vertex $j$ to vertex $k$, the entry in the upper triangular part of the adjacency matrix transforms as $1\rightarrow\exp(-i[\theta_{j}-\theta_{k}])$, with the corresponding complex conjugate value in the lower triangular part. The diagonal elements remain as entries of 0.
In the numerical examples shown below we obtain the phases from simulations using the Kuramoto model, described in the previous section. For any time point during the simulation, we use the oscillator phases to define $\Phi$, then we diagonalize $\Phi^{-1}A\Phi$ to obtain the eigenvectors of the emergent states—now with complex coefficients. We project these coefficients to those for an effective two-state system for each QL bit, as we now describe for the example of a single QL bit. The method is generalized in the obvious way for a system comprising any number of QL bits, as shown in Sec. V.C.
@@ IV.2 The effective 2-state model for QL bit emergent states
An arbitrary emergent state, $W_{i}$, of the QL bit graph will be written in terms of this basis as:
$W_{i}=c_{1}u_{1}+\dots+c_{n}u_{n}+c_{n+1}x_{1}+\dots+c_{n+k}x_{k}$ (14)
where the $c_{l}$ are complex coefficients. Now define
@@
$\alpha_{i}=\langle J_{a1},W_{i}\rangle$ (17)
$\beta_{i}=\langle J_{a2},W_{i}\rangle$ (18)
where $i$ labels the state of the QL bit, ordered by eigenvalue. We thus obtain the effective states of the QL bit as
$W_{i}^{2\times 2}=\alpha_{i}|a_{1}\rangle+\beta_{i}|a_{2}\rangle.$ (19)
Some examples of coefficients for the effective two-states of a QL bit are provided in Fig. 2. Each panel shows a vector representation of the complex coefficients of $W_{i}^{2\times 2}$ corresponding to the highest eigenvalue, $i=0$ ($\alpha_{0}$ is violet red, $\beta_{0}$ is light pink) and the second highest eigenvalue, $i=1$ ($\alpha_{1}$ is dark blue, $\beta_{1}$ is light blue). Fig. 2a shows the eigenvectors for partially synchronized oscillators in one QL bit, derived from phases early in the dynamics of a phase oscillator system with initially random phases. The simulation used the same parameters as detailed later for the simulations of synchronization on the product graph of two QL bits. These coefficients contrast to those obtained for the synchronized oscillators in one QL bit, where phases were obtained later in the simulation once the oscillators has synchronized, Fig. 2b.
@@ IV.3 Phase in the states
Figure 2: Vector representation of the complex coefficients of $W_{i}^{2\times 2}$ corresponding to the highest eigenvalue , $i=0$ ($\alpha_{0}$ is violet red, $\beta_{0}$ is light pink) and the second highest eigenvalue, $i=1$ ($\alpha_{1}$ is dark blue, $\beta_{1}$ is light blue). (a) Examples of systems comprising partially synchronized oscillators. (b) Examples of systems where the oscillators are synchronized.
For a synchronized oscillator system on the QL bit graph with all edges set to 1 in the adjacency matrix, we find that $\alpha_{0}=\beta_{0}\approx 1/\sqrt{2}$ and $\alpha_{1}=-\beta_{1}\approx 1/\sqrt{2}$ (we use the convention that $\beta_{1}$ takes the negative sign). More generally, we can bias the edges that connect the two subgraphs by giving them values (in the adjacency matrix) of any value on the unit circle in the complex plane. Recall that the special case of negative-signed edges connecting the subgraphs can be interpreted as phases of the two subgraphs being perfectly out of phase[1]. The resulting eigenvaectors can be evaluated by considering the effective adjacency matrix for the QL bit graph comprising two $d$-regular subgraphs coupled by edges given the bias $e^{i\alpha}$:
$A_{2\times 2}=\begin{bmatrix}d&e^{i\gamma}\\ e^{-i\gamma}&d\end{bmatrix},$ (20)
which gives an eigenvector for the emergent state $(e^{i\gamma},1)$ and for the state associated with the next eigenvalue $(-e^{i\gamma},1)$. These states are obtained by continuous rotation of the connecting edge bias. Note that we have neglected an overall phase of these states. When that phase is accounted for, then we see that the emergent state is any vector in $\mathbb{C}^{2}$.
By rotating the edge bias, we therefore generate states (when the system is perfectly synchronized) that emulate, for example, the states of a system of spins or polarization states of photons. The notable emergent states isomorphic to polarization states of light are: (i) the linear combinations $a_{1}\pm a_{2}$, when edges take values ±$\pm 1$ respectively, which are comparable to linear combinations of vertical and horizontal polarization; (ii) the states $a_{1}\pm ia_{2}$, when edges take values of ±$\pm i$, analogous to right- and left-circularly polarized light.
@@ IV.4 Synchronization of quantum bits
$i\hbar\dot{U}_{i}U^{\dagger}_{i}=H_{i}-\frac{iK}{2N}\sum_{j=1}^{N}a_{ij}(U_{j}U^{\dagger}_{i}-U_{i}U^{\dagger}_{j}).$ (21)
@@
It seems that we can emulate this synchronization of quantum bits by building a system of QL bits, interacting via edges connecting vertices of one QL bit to another. For consistency with the Lohe model, we would require each QL bit to comprise strongly synchronized oscillators, and to be synchronized. The emergent state of each QL bit gives its frequency. We can assign a random initial ‘phase’ as the real 4-vector vector in $\mathbb{S}^{3}$ that defines the orientation of the eigenvector of the synchronized emergent state. The key point is that this vector, associated to $|\psi_{i}\rangle$ in Eq. 6 of Ref. [43], must represent the emergent state vector as an element that sweeps through values in $\mathbb{C}^{2}$, and we have established this to be the case.
The system of $N$ QL bits would be combined into a new graph $G=G_{A}\oplus G_{B}\oplus\dots\oplus G_{N}$. The subgraphs are connected by edges, weighted so that the inter-QL bit coupling is weaker than the coupling between vertices within a QL bit. The graphs $G_{A},G_{B},\dots$ are the vertices of the Lohe model. This system of oscillators, propagated according to the Kuramoto model, should mimic the Lohe model. To compare the results we should project out the effective 2-state emergent states of each QL bit to construct the vectors $\mathbf{x}_{i}$. We do not report any associated numerical calculations here, but we will refer to this model in a later section of the paper. By proposing how a model comprising a system of coupled classical phase oscillators can predict the same results as a system of coupled quantum bits emphasizes how the QL model succeeds in various ways to produce state spaces very much like those enjoyed by quantum systems.
@@ V.1 Methods
Here we study a phase oscillator network model with a coupling structure templated by the Cartesian product of two QL bit graphs $G_{A}\Box G_{B}$. The phase oscillators are indicated by vertices of the graph, and are assigned a phase that biases the edges, as we describe below.
@@
In the numerical simulations, each subgraph comprises $n_{0}$ vertices, giving a total of $N=2n_{0}\times 2n_{0}$ vertices in the product graph. For the calculations, each subgraph $G_{a1}$, $G_{a2}$, $G_{b1}$, $G_{b2}$ comprises 20 vertices and is $d$-regular with $d=15$. Edges between subgraphs were added randomly, with probability 0.2. We calculate the full Cartesian product, yielding $2d$-regular subgraphs in $G_{A}\Box G_{B}$, with an eigenvalue of 30 for the emergent state when the subgraphs are uncoupled (that is, neglecting the eigenvalue shift labeled $\Delta$ in Ref. [2]). We expect four emergent states, two with an eigenvalue of 30, one at $30+\Delta$, and one at $30-\Delta$.
@@
The graph provides the coupling map for a system of $N$ phase oscillators that we model using the Kuramoto model, Eq. 10. Couplings between vertices are weighted uniformly by $K/N$, with $K$ chosen to be 250 for the first set of calculations. Other choices of $K$ are reported in the subsequent results. Vertices are assigned a mean frequency of 100 in the same arbitrary units as $K$, and offset by frequencies randomly selected from a normal distribution with standard deviation 1 (in the same arbitrary units). Initial phases for the oscillators are randomly assigned from a von Mises distribution with a very large circular standard deviation, meaning the distribution is similar to a uniform distribution, or 0.001, meaning the distribution is extremely narrow. Results are averaged over either 100 or 500 realizations of this set-up. See ref [40] for some background and examples that might be helpful. In the following subsection we discuss the Kuramoto model to provide context and justification for the choice of these parameters.
@@
The main motivation for the averaging over realizations of the initial phases is to account for uncertainty in preparation of the initial state. Therefore, we use the term ‘ensemble’ to mean an average over initial conditions of oscillator phase. The initial phase for any oscillator refers to its natural frequency together with its phase offset, $\epsilon_{i}$ and $\phi_{i}$, respectively. This setup for the calculations means that the dynamics are not strictly deterministic, because we do not know the initial conditions precisely. We also checked calculations where we, additionally, average over an ensemble of graphs. The subgraphs are $d$-regular (with none of the edges randomly deleted like we did in prior work[1]), so the ensemble comprises graphs with different configurations of edges connecting the subgraphs.
@@
During the evolution of the system, the phase of each oscillator $\theta_{i}$ evolves nonlinearly, as discussed in the previous section. At selected time points we transform the adjacency matrix of the graph according to Eqs. 12, 13, which introduces phase differences by weighting the edges between vertices $j$ and $k$ by $e^{i(\theta_{k}-\theta_{j})}$. We average the spectrum of the transformed adjacency matrix over all iterations of the simulation to give ensemble spectra. The spectra were computed for various time points, and therefore order parameters, as the system synchronized. Note that although the phases of the initial distribution of oscillators are randomly distributed, the phase differences contain embedded correlations—there are $N$ phases, but much more than $N$ phase differences. However, the phase differences are uncorrelated from one simulation to another, so mixedness comes from an essential ensemble averaging over simulations. In prior work mixedness as a function of synchronization was studied in some detail[40].
V.2 Parameters and State preparation
The evolution of the oscillator system, though sensitive to the specified initial conditions, is entirely deterministic. However, we use a particular averaging over the initial conditions that reflects our hypothesis that we cannot know the initial phases of the oscillators, only the variance of the distribution. These initial phase conditions are decided by a probabilistic state preparation. To account for such a state preparation, in the numerical studies we take an ‘ensemble’ average over initial conditions of the phase oscillators which is intended to capture the fact that we cannot know precisely the phases. It is not an average over a set of distinct QL systems.
The parameters in the calculations are selected using well-known properties of the Kuramoto model, Eq. 10. Whether or not a system synchronizes depends on the size of the coupling strength $K$ compared to the spread of frequencies $\nu_{i}$. This principle is clearly illustrated by a system of only two phase oscillators, see Sec. 8.6 of Ref. [46]. The phase portrait (i.e. $\theta_{1}$ is plotted versus $\theta_{2}$ on the torus) shows stable phase-locked solutions when $|\nu_{1}-\nu_{2}|<2K$. That is, the phase oscillators synchronize. Whereas, when $|\nu_{1}-\nu_{2}|>2K$ the flow resembles that of uncoupled oscillators (the coupling is insufficient to enable them to synchronize). The $|\nu_{1}-\nu_{2}|=2K$ is a kind of dividing line, a saddle point bifurcation.
When there are many oscillators, the situation is more complex because we can find situations where some oscillators in the system synchronize with each other and others do not. That outcome will depend on the distribution of frequencies as well as the way the oscillators are coupled (i.e. the graph that defines the coupling matrix $[a_{ij}]$). Nevertheless, a rule of thumb for a system with all-to-all coupling is that when $K$ in Eq. 10 sufficiently exceeds the standard deviation of oscillator frequencies ($\sigma_{\nu}$), the system will synchronize. For examples see Ref. [25, 26]. The network studied here does not have all-to-all coupling, so we tested various couplings to find parameters where the system of phase oscillators will synchronize or will not maintain synchronization.
In more detail, whether or not a system of oscillators will perfectly synchronize depends not only on $K$ and $\sigma_{\nu}$, but also on the initial conditions. That is, the specific assignment of frequencies and initial phases to each oscillator in the network. For insight into this statement see, for example, Refs. [10, 11, 9, 47, 48]. This sensitivity of the system to initial conditions emphasizes the nonlinearity of the network.
V.3 Observation of a State
@@
According to Eq. 12 we are applying a unitary transformation to the adjacency matrix that is decided by the oscillators phases at any time. Therefore the spectrum does not change with order parameter, only the state. The density matrix for the relevant emergent state is calculated in the basis produced by the effective two-state model for each QL bit. That is, we statistically average over the pure states:
$W_{i}^{4\times 4}(t)=c^{i}_{1}(t)|a_{1}\rangle|b_{1}\rangle+c^{i}_{2}(t)|a_{2}\rangle|b_{1}\rangle+c^{i}_{3}(t)|a_{1}\rangle|b_{2}\rangle c^{i}_{4}(t)|a_{2}\rangle|b_{2}\rangle,$ (25)
for each time point in a simulation. Here the $c_{k}(t)$ are complex coefficients obtained from the projections analogous to those outlined in Sec. IV.B. We obtain the coefficients as follows. We form the Cartesian product of two QL bit graphs, say $G_{A}$ and $G_{B}$, each comprising two subgraphs with $n_{0}$ vertices. From the product, $G_{A}\Box G_{B}$, we end up with four subgraphs (each with $n_{0}^{2}$ vertices). The subgraphs are connected with a certain structure of edges that define the correlations. The vertices in each subgraph are labeled by all possible $n_{0}^{2}$ pairs of vertex indices, where one index is taken from a subgraph of $G_{A}$ and the other from a subgraph of $G_{B}$. We project the emergent states of the entire graph to effective states in a chosen basis. Here we choose the natural product basis associated directly with the subgraphs of $G_{A}\Box G_{B}$: $|a_{1}\rangle|b_{1}\rangle$, $|a_{2}\rangle|b_{1}\rangle$, $|a_{1}\rangle|b_{2}\rangle$, $|a_{2}\rangle|b_{2}\rangle$. That basis is selected by defining the vectors
$J_{a1b1}=\{\underbrace{1,1,\dots,1}_{n_{0}^{2}\text{ times}}\underbrace{0,0,\dots,0}_{3n_{0}^{2}\text{ times}}\}/\sqrt{N}$ (26)
$J_{a2b1}=\{\underbrace{0,0,\dots,0}_{n_{0}^{2}\text{ times}}\underbrace{1,1,\dots,1}_{n_{0}^{2}\text{ times}}\underbrace{0,0,\dots,0}_{2n_{0}^{2}\text{ times}}\}/\sqrt{N}$ (27)
etc. (28)
where it is assumed for simplicity that the ordering of the vertices in the graph product partitions the subgraphs in sequence (which may not be the case in numerical work). Then we obtain coefficients for the effective four states of the QL bit product in terms of the product basis naturally arising from the subgraphs:
$c^{i}_{1}=\langle J_{a1b1},W_{i}\rangle$ (29)
$c^{i}_{2}=\langle J_{a2b1},W_{i}\rangle$ (30)
etc. (31)
where $i$ labels the state of the QL bit, ordered by eigenvalue.
From the average over state preparations of $W_{i}^{4\times 4}(t)$ we determine the density matrix $\rho(t)$ associated with the emergent state $i$ in the product basis, with entries
$\rho_{mn}(t)=\sum_{\text{prep}}c_{m}(t)c^{*}_{n}(t)$ (32)
where $\sum_{\text{prep}}$ means an average over state preparations. We report $\mathrm{Tr}(\rho)^{2}$ as a simple measure of mixedness of the state. Notice that the complex phases cancel to unity in the diagonal elements of $\rho_{mn}(t)$. On the other hand, the complex phases are retained in the off-diagonal elements and unless the system is synchronized, they interfere in the average and can thereby diminish or erase the off-diagonal elements, introducing state mixedness.
The vectors $J_{a1b1}$ and so on define the measurement basis for the QL state. Here we set them up so that the basis is the product basis associated directly with the subgraphs of $G_{A}\Box G_{B}$. We could, however, choose any other basis. For instance, we could select the Bell basis using
$J_{a1b1\pm a2b2}=\{\underbrace{1,1,\dots,1}_{n_{0}^{2}\text{ times}}\underbrace{0,0,\dots,0}_{2n_{0}^{2}\text{ times}}\underbrace{1,1,\dots,1}_{n_{0}^{2}\text{ times}}\}/\sqrt{N}$ (33)
$J_{a1b2\pm a2b1}=\{\underbrace{0,0,\dots,0}_{n_{0}^{2}\text{ times}}\underbrace{\pm 1,\pm 1,\dots,\pm 1}_{n_{0}^{2}\text{ times}}\underbrace{1,1,\dots,1}_{n_{0}^{2}\text{ times}}\underbrace{0,0,\dots,0}_{n_{0}^{2}\text{ times}}\}/\sqrt{N}.$ (34)
A preferred basis could be set up by suitable weighting of the edges.
V.4 Synchronization